The Brenmar Sales Company had a gross profit margin​ (gross profits divided by ​sales) of 25 percent and sales of $ 8.1 million last year.  75 percent of the​ firm's sales are on​ credit, and the remainder are cash sales. ​ Brenmar's current assets equal $ 1.3 ​million, its current liabilities equal $ 301300​, and it has $ 102900 in cash plus marketable securities. a. If​ Brenmar's accounts receivable equal $ 562400​, what is its average collection​ period? b. If Brenmar reduces its average collection period to 15 ​days, what will be its new level of accounts​ receivable? c.  ​Brenmar's inventory turnover ratio is 9.9 times. What is the level of​ Brenmar's inventories? a.  If​ Brenmar's accounts receivable equal $ 562400​, what is its average collection​ period? The​ company's average collection period will be nothing days.  ​(Round to two decimal​ places.)
Answer #1

a)

Credit sales = 8,100,000 * 0.75 = 6,075,000

Receivables turnover ratio = Credit sales / average receivables

Receivables turnover ratio = 6,075,000 / 562,400

Receivables turnover ratio = 10.80192

Average collection period = 365 / 10.80192

Average collection period = 33.79 days

b)

15 = 365 / Receivables turnover

Receivables turnover = 24.3333

24.3333 = 6,075,000 / average receivables

New level of inventory = $249,657.88

c)

Gross profit = 8,100,000 * 0.25 = 2,025,000

COGS = 8,100,000 - 2,025,000

COGS = 6,075,000

Inventory turnover ratio = COGS / inventory

9.9 = 6,075,000 / inventory

inventory = 613,636.36